# ICAS: Detecting Training Data from Autoregressive Image Generative Models

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ACM Multimedia 2025; October 27-31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia (MM ’25), Octobe 27–31, 2025, Dublin, Irelandisbn: 978-1-4503-XXXX-X/2018/06submissionid: 174ccs: Security and privacy Privacy protectionsccs: Computing methodologies Computer vision
@@ 4. Experiments
Evaluation Metrics. Following previous researches (Li et al., 2024a; Shi et al., 2023), to evaluate our method, we adopt the widely used metrics, including Area-Under-the-ROC-curve (AUROC), and the True Positive Rate when the False Positive Rate is $5\%$ (TPR@$5\%$), and attack success rate (ASR). For ASR, we randomly select a subset of both member and non-member samples, comprising $20\%$ of the total, and compute an optimal threshold from it. Then we apply the threshold to evaluate the ASR for the remaining samples.
Baselines. Since the autoregressive generation paradigm in visual autoregressive models is similar to the text generation paradigm, we adapt some advanced training data detection methods for large language models as our baseline. Specifically, we adapt the $Loss$ attack (Yeom et al., 2018), $Min$-$k\%$ (Shi et al., 2023), $Min$-$k\%$++ (Zhang et al., 2024b) and Rényi entropy score (Li et al., 2024c) in our experiments. The implementation detail for baselines is provided in Appendix A.
@@ 4.5. MI Linear Scaling Law
(17) $\mathrm{AUROC}_{\text{NAR}}$ $=0.143P+0.517$ $r$ $=0.998$
$\mathrm{AUROC}_{\text{LlamaGen-CCA}}$ $=0.157P+0.511$ $r$ $=0.999,$
@@ A.3. Min-k%++
The $Min$-$k\%$ method (Zhang et al., 2024b) is an upgrade from $Min$-$k\%$. It considers the expected probability of all tokens within the vocabulary. The token score is:
@@ C.2. Baseline methods
$\mathrm{AUROC}_{Loss}$ $=0.1538P+0.4614$ $r=0.9946$
$\mathrm{AUROC}_{Min-k\%}$ $=0.1934P+0.4715$ $r=0.9995$
$\mathrm{AUROC}_{Min-k\%++}$ $=0.1571P+0.4547$ $r=0.9904$

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ACM Multimedia 2025; October 27-31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), Octobe 27–31, 2025, Dublin, Irelandisbn: 978-1-4503-XXXX-X/2018/06submissionid: 174ccs: Security and privacy Privacy protectionsccs: Computing methodologies Computer vision
@@ 4. Experiments
[…]
Baselines. Since the autoregressive generation paradigm in visual autoregressive models is similar to the text generation paradigm, we adapt some advanced training data detection methods for large language models as our baseline. Specifically, we adapt the $Loss$ attack (Yeom et al., 2018), $Min$-$k\%$ (Shi et al., 2023), MinMin-$k\%$++ (Zhang et al., 2024b) and Rényi entropy score (Li et al., 2024c) in our experiments. The implementation detail for baselines is provided in Appendix A.
@@ 4.5. MI Linear Scaling Law
[…]
@@ A.3. Min-k%++
The $Min$-$k\%$ method (Zhang et al., 2024b) is an upgrade from MinMin-$k\%$. It considers the expected probability of all tokens within the vocabulary. The token score is:
@@ C.2. Baseline methods
[…]